solve the system
3x+4y=40
x-2y=0

Answer by scoop(27)

3x+4y=40
x-2y=0
I would use the elimination method and get rid of the y variable:
3x+4y=40
(2)x-2y=0 leaving you with 2x-4y=0
then you simply add equations
3x+4y=40
2x-4y=0
5x=40 divide both sides by 5
x=8
to check...substitute 8 in expression for x
3(8)+4y=40
24+4y=40
4y=40-24
4y=16
y=4
24+16=40
8-8=0
you are correct!!!